OpenStudy (anonymous):

two numbers total 33 and have a difference of 5. Find the two numbers.

OpenStudy (anonymous):

The larger number is ___ and the smaller number is__

OpenStudy (hunus):

Well, let's call the two numbers x and y. One of them must be greater than the other if their difference is 5 so let's let the bigger one be x. Then we know that x + y is equal to 33 and the difference is positive so it's the bigger one minus the smaller one or x - y so that x - y = 5 So now we have x + y = 33 and x - y = 5 Do you know where to go from here?

OpenStudy (anonymous):

no

OpenStudy (hunus):

Do you know how to manipulate equations like if a - b = 10 then a = 10 + b?

OpenStudy (hunus):

Well each operation in arithmetic has it's opposite. The opposite of addition is subtraction, the opposite of multiplication is division; so when you have things on different sides of the equal sign you can put them on the other side by doing the opposite operation on each side. For example, a + b = 12 if we want to move b to the opposite side we put the opposite of b on each side, or -b a + b - b = 12 - b but b - b = 0 so we're left with a = 12 - b Which makes sense if you think if a and b add together to get 12, then 12 minus b should be a 7 + 5 = 12 12 - 7 = 5 In the above problem we have x + y = 33 and x - y = 5 so that x is 5 larger than y or x = y + 5 Now we can just replace the x in x + y = 33 with y + 5 because they are the same thing so now we have y + 5 + y = 33 or 2y + 5 = 33 Do you know what you could do from here?